OpenStudy (anonymous):

If 21.7 mL Na, with a density of 0.97 g/mL, react with excess oxygen gas, how many moles of sodium oxide can be formed? unbalanced equation: Na + O2 “yields”/ Na2O 0.458 mol Na2O 0.487 mol Na2O 1.83 mol Na2O 1.95 mol Na2O (I know this is chem but no one is alive in that section so I really, really am just desperate and need someone to help me out. Please)

OpenStudy (cwrw238):

use the equation 2Na + O2 = Na2O 2 moles of sodium give one mole of sodium oxise

OpenStudy (cwrw238):

sorry i havent balanced the equation its 4 Na + O2 = 2Na2 O so 2 moles of sodium give 1 mole of sodium oxide

OpenStudy (anonymous):

Yeah I did that. And for the stoichiometry part, I did: 21.7 mL Na * 1 mol Na/22.989 g Na * 2 mol O/4 mol Na * 0.97/1 mol

OpenStudy (anonymous):

But I'm not sure what the answer would be. I'm not even sure if my work is correct or not

OpenStudy (anonymous):

Do you know what the answer is @cwrw238 ?

OpenStudy (anonymous):

I think it's A

OpenStudy (cwrw238):

long time since i've done this stuff first i think you must convert 21.7 mls of Na to mole 21.7 mls of 0.97 g / ml = 21.7 * 0.97 g = 21.31 g sodium

OpenStudy (cwrw238):

1 mole of sodium = 23 g right?

OpenStudy (anonymous):

yes

OpenStudy (cwrw238):

so thats 0.9265 moles of sodium so if 2 moles gives 1 mole of sodium oxide then we will get 0.9265 / 2 = 0.4633 moles of the oxide

OpenStudy (cwrw238):

which is close to A..

OpenStudy (cwrw238):

i cant see where i;ve made any mistake...

OpenStudy (cwrw238):

is 23 g right for 1 mole of Na?

OpenStudy (cwrw238):

22.99 is given on website - that will make very little difference

OpenStudy (cwrw238):

oh - can see my mistake 21.7 * 0.97 = 21.049 g = 0.916 moles giving 0.916 = 0.458

OpenStudy (cwrw238):

yes its A ok/

OpenStudy (cwrw238):

I must have pressed wrong key on my calculator lol

OpenStudy (cwrw238):

0.916 / 2 = 0.458

OpenStudy (cwrw238):

do you follow my method ok?

OpenStudy (anonymous):

oh hey sorry for replying so late. And yes, thank you

OpenStudy (cwrw238):

yw